OpenStudy (anonymous):

find the product and simplify it possible. 7s^5y^3/21s^5y^5.y^6

OpenStudy (saifoo.khan):

\[\frac{7s^5y^3}{21s^5y^5} \times y^6\]Like that?

OpenStudy (anonymous):

0.33333y^14

OpenStudy (anonymous):

is that the final answer?

OpenStudy (saifoo.khan):

Is that what i wrote?

OpenStudy (anonymous):

O_o no?

OpenStudy (saifoo.khan):

THen?

OpenStudy (anonymous):

XD -confused-

OpenStudy (saifoo.khan):

why? LOL.

OpenStudy (anonymous):

no idea..just been fustrated the whole night. XD

OpenStudy (saifoo.khan):

lol, how does the question looks like? does it looks like how i wrote it?

OpenStudy (saifoo.khan):

Can u draw it or post a photo or something?

OpenStudy (saifoo.khan):

that's the same as i wrote. LOL just there is no negative sign with u didnt wrote either.

OpenStudy (anonymous):

ha?

OpenStudy (saifoo.khan):

\[-\frac{7s^5y^3}{21s^5y^5} \times y^6\]

OpenStudy (anonymous):

maybe there is a chat lag, I didn't see it before

OpenStudy (saifoo.khan):

\[-\frac{y^4}{3}\]

OpenStudy (anonymous):

I got everything right but the 4...

OpenStudy (anonymous):

but thanks

OpenStudy (saifoo.khan):

No problem.